Graph the line with slope -2 passing through the point (-1,-1).
Found 3 solutions by ikleyn, Alan3354, math_tutor2020:
Answer by ikleyn(53937)

.

Use graph paper, a pen and a straightedge.



Answer by Alan3354(69443)

Graph the line with slope -2 passing through the point (-1,-1).
----------------------
OK, I did that. Now what?
=====================
If you want to do it too, on graph paper plot the given point.
A slope of -2 means each change of 1 in x results in a change of -2 in y.
Plot a 2nd point by add 1 to x, and -2 to y. ---> the point (-1+1,-1-2).
Plot that point.
Draw a straight line thru the 2 points.

Answer by math_tutor2020(3838)


slope = -2 = -2/1

slope = rise/run = -2/1
rise = -2 = go down 2 units
run = 1 = go to the right 1 unit

Start at (-1,-1)
Move down 2 and to the right 1 to arrive at (0, -3)

This is what it should look like

The stuff in red is to be erased later. It's meant to be a visual tool to see what's going on with the slope.
Though your teacher may want you to visually describe these actions. Ask him/her for clarification.

I used GeoGebra to make the graph. Specifically I used the desktop version, but the web app and phone app are equally useful of course.
Desmos is another option I recommend as well.
Both apps mentioned are free.